OpenStudy (anonymous):

How do I find the exact value of of sec (-5pi/3)??

hero (hero):

replace pi with 180,then continue solving

hero (hero):

sec(-5(180)/3) = sec(-5(60)) = sec(-300) = 1/cos(-300) = 1/.5 = 2

OpenStudy (anonymous):

Thanks ton!

OpenStudy (anonymous):

How did you get to the 1/.5 step?

OpenStudy (anonymous):

Oh nevermind

OpenStudy (unklerhaukus):

cos (-300) = cos (60)

OpenStudy (anonymous):

Ok that makes sense thanks